21. 4.2 mL of isoamyl alcohol was reacted with 3.5 mL of acetyl chloride and 2.7 grams of isoamyl acetate was recovered, what is the yield? Include calculations for density conversions the limiting reagent, theoretical yield and yield answer:
Answer #1

First,

identify the theoreticla amount of isoamyl acetate

so

mass of Isoamyl alcohol

D = 0.810 g/mL

mass = D*V = 0.81*4.2 = 3.402 g of Isoamyl Alcohol

mol = mass/MW = 3.402/88.148 = 0.03859 mol of Isamyl alcohol

then

mass of Acetyl chloride = D*V

D = 1.1 g/mL

mass = D*V = 1.1*3.5 = 3.85 g of acetly chloride

now

mol = mass/MW = 3.85/78.49 = 0.049

ratio is 1:1

so

0.03859 : 0.049

0.03859 mol will be formed

0.03859 mol of isoamyl acetate

mass = mol*MW = (0.03859 )*130.19 = 5.0240 g should form at 100%

so

%yield = real / theoretical * 100%

% yield = 2.7/5.0240*100 = 53.742 %